OpenStudy (anonymous):

I need someone help to get the right answer? -4x-12.85=25.25+9x

OpenStudy (amistre64):

combine like terms

OpenStudy (amistre64):

-12.85 = 9x -25.25 4x ------------ -38.10 = 13x x = 18.1/13

OpenStudy (amistre64):

reduce as wanted :)

OpenStudy (anonymous):

thank you amistre64 I really appreciate ur help.

OpenStudy (amistre64):

youre welcome :)

OpenStudy (gw2011):

I believe the answer should be x = -38.1/13

OpenStudy (amistre64):

its hard to type in the dark :)

OpenStudy (amistre64):

but youre correct :)